# Trustworthy Semantic-Enabled 6G Communication: A Task-oriented and Privacy-preserving Perspective

### v1

@@ Abstract
Trustworthy task-oriented semantic communication (ToSC) emerges as an innovative approach in the 6G landscape, characterized by the transmission of only vital information that is directly pertinent to a specific task. While ToSC offers an efficient mode of communication, it concurrently raises concerns regarding privacy, as sophisticated adversaries might possess the capability to reconstruct the original data from the transmitted features. This article provides an in-depth analysis of privacy-preserving strategies specifically designed for ToSC relying on deep neural network-based joint source and channel coding (DeepJSCC). The study encompasses a detailed comparative assessment of trustworthy feature perturbation methods such as differential privacy and encryption, alongside intrinsic security incorporation approaches like adversarial learning to train the JSCC and learning-based vector quantization (LBVQ). This comparative analysis underscores the integration of advanced explainable learning algorithms into communication systems, positing a new benchmark for privacy standards in the forthcoming 6G era.
@@ I Introduction
The sixth generation (6G) communication represents the next frontier in communication technology, promising significant advancements over the current 5th generation (5G) networks. 6G networks are projected to achieve data speeds in the range of terabits per second, a significant leap from the gigabit speeds of 5G [1]. The integration of artificial intelligence (AI) in 6G networks is anticipated to be more profound, with AI algorithms playing a crucial role in network management and user-centric services. Additionally, 6G is set to revolutionize industries by enabling new business models and services. The network’s reliability, trustworthiness and timeliness will be critical in the scenario of massive mobile users with real-time response requirements [2]. Despite the promising prospects, the practical implementation of 6G continues to encounter numerous unprecedented hurdles, particularly for burst communications [3]. ToSC, emerging as a promising paradigm, is primarily characterized by its selective transmission of information [4]. This approach has garnered considerable attention [5], chiefly due to its proficiency in enhancing efficiency and reducing latency through the minimization of data transmission volume [6]. Furthermore, ToSC exhibits the capacity to offer a more customized and efficient user experience. The targeted and efficient nature of ToSC, therefore, not only aligns with the technological advancements envisaged in 6G communications but also caters to the nuanced demands of modern digital applications, ensuring a seamless and user-oriented interaction.
@@
To alleviate the misunderstandings and incorrect interpretations, ToSC can increase information clarity, relevance, transparency, credibility, and verifiability by concentrating on task-relevant information, implementing well-designed mechanisms of channel coding and feedback. Although this strategy inherently provides a certain level of privacy since selective data transmission eliminates the unnecessary data sharing, the information conveyed may still be vulnerable [7]. If intercepted, even these task-relevant data bits could reveal personal or sensitive information [8]. This risk is heightened by the advanced capabilities in machine learning and data analysis, which might enable adversaries to extract significant insights from minimal data. This work discusses ongoing research aimed at developing privacy-preserving methods specifically for ToSC relying on deep neural network-based joint source and channel coding (DeepJSCC). The task at hand is devising techniques that can be seamlessly incorporated into semantic communication without hindering their efficacy and efficiency. This work meticulously evaluates and contrasts feature perturbation methodologies, such as differential privacy and encryption techniques, with intrinsic security incorporation approaches like adversarial learning to the JSCC and learning-based vector quantization (LBVQ). This comprehensive comparative analysis highlights the potential for integrating sophisticated learning algorithms into contemporary communication systems, thereby establishing a new paradigm for privacy standards in the forthcoming 6G era.
@@ II Shift From Task-Agnostic Communications to ToSC and Privacy Challenges
Within the traditional framework of source-channel separation, the identification, representation, and transmission of information are rigorously addressed by rate-distortion theory and channel coding theory respectively. This paradigm, which prioritizes reconstruction-oriented compression and task-agnostic communication, has underpinned several iterations of digital communication systems. However, the advent of machine-to-machine communications and human-machine interactions necessitates a reassessment of this paradigm, considering that exact reconstructions are often of secondary importance from a machine’s perspective. Notably, task-specific descriptors, derived via machine learning algorithms from latent feature spaces, are substantially more concise than their counterparts used for reconstruction purposes. Furthermore, communication systems trained end-to-end can significantly surpass those designed based on source-channel separation, across various performance metrics [4].
@@
In ToSC systems, data transmission is meticulously tailored to align with the receiver’s requirements. As delineated in Fig. 1, the ToSC transceivers differ from traditional transceivers. The ToSC system comprises two main components: the “DeepJSCC” and the “Task inference module”. The DeepJSCC contains a JSCC encoder that plays the role of task-related information extraction, compression, and protection (against both channel noise and adverseries’ attacks). The output of this passes through a channel and arrives at the task inference module. This module is outlined with examples of tasks such as classification, object detection, information retrieval, and content generation. Inside, it displays a multi-layered network topology that indicates intricate processing is occurring to perform the given tasks.
@@
However, alongside these advantages, ToSC introduces significant privacy concerns. The main issue stems from the nature of the information being transmitted. Although ToSC systems transmit only task-specific information, this data can be sensitive and vulnerable to privacy infringements. Given the unpredictability of adversaries’ objectives, it is imperative to devise a comprehensive and efficacious strategy for safeguarding a spectrum of private data. For instance, in the transmission of facial imagery, adversaries may undertake diverse strategies to extract personal attributes such as gender or skin tone, or alternatively, engage in facial recognition. In our research, we postulate the quality of image reconstruction—evaluated by metrics such as peak signal-to-noise ratio (PSNR)—as a quantifiable proxy for privacy preservation. Our rationale is predicated on the supposition that by amplifying the perturbation of data reconstructed by potential intruders, we indirectly shield various facets of personal information. This approach resonates with the principles of perturbative privacy preservation, notably exemplified by differential privacy paradigms. By escalating the degree of distortion, we can significantly diminish the likelihood or amplify the challenge for adversaries in gleaning sensitive information, thereby strengthening the robustness of privacy protection.
@@ III Privacy Preservation Methods for ToSC
Most ToSC systems are based on DeepJSCC architectures and employ an end-to-end training methodology to extract high-dimensional task-related channel-robust features for transmission. This approach ensures a coherent and automatically optimized process from data input to the final task output. Many traditional privacy protection techniques, such as $k$-anonymity, $l$-diversity, and $t$-closeness, are often designed to protect user privacy by modifying data and are not suitable for working with complex or high-dimensional data [8, 9], as the process of making records indistinguishable can lead to significant data loss or impracticality in datasets with numerous attributes. A thorough analysis of privacy-preserving strategies specifically designed for ToSC is provided, along with a detailed comparative assessment of feature perturbation methods and intrinsic security incorporation approaches.
@@ III-A Feature Perturbation Methods
Conventional approaches to AI inference privacy protection concentrate primarily on the safeguarding of original data, typically through direct preprocessing under the assumption that the entire AI model is in the possession of a potentially untrustworthy third party. However, in DeepJSCC-based ToSC, the DeepJSCC encoder (partial of the AI models) belongs to the transmitter, which plays the role of task-related information extraction, compression, and protection (against both channel noise and adverseries’ attacks). The employment of DeepJSCC makes conventional privacy and security techniques applied in-between source and channel coding difficult in DeepJSCC. Applying these techniques in front of DeepJSCC will prevent the DeepJSCC to extract the task-related and channel-robust features, thus the only feasible is to process the output of DeepJSCC. Therefore, traditional data perturbations (e.g., noise addition, encryption) should be replaced by the feature perturbations on the DeepJSCC output. Next, we discuss two typical feature perturbation methods for privacy preservation.
  • Differential privacy offers a mathematical framework for quantifying privacy loss. It provides strong privacy guarantees by adding noise to the DeepJSCC output. This approach is widely adopted due to its robustness and the theoretical guarantees. The addition of noise to protect individual data points can diminish the utility of data, especially in scenarios where precise data (i.e., semantic information) is used for task inference, as in ToSC. Too much noise can well protect privacy, however, it will inherently affect the task inference performance, that is the utility of the transmitted data.
  • Encryption serves as a fundamental method for securing the output of DeepJSCC, by converting it into a format that is not readily intelligible or accessible without the appropriate decryption key. However, implementing encryption and decryption processes can significantly decelerate both the training and inference phases of ToSC. Additionally, in the absence of channel coding, encryption becomes vulnerable and may compromise the integrity of decryption. To mitigate these challenges, only methods such as simple encryption or error-robust encryption techniques are recommended for ToSC. An example of such a technique is the key-guided feature shuffling, as discussed in the recent literature [10]. Besides, encryption-based privacy-preserving methods are facing an additional key-sharing burden and the challenge of key leakage.
@@ III-B Intrinsic Security Incorporation Strategies
It is key to design and train DeepJSCC with intrinsic security. Adversarial learning, as a commonly-used technique, involves training a model in a competitive setting where two models, usually called the generator and the discriminator, are pitted against each other. The generator aims to produce data that is indistinguishable from real data, while the discriminator’s goal is to accurately distinguish between the generator’s fake data and real data. This process helps to improve the performance in both models. Adversarial learning plays a crucial role in privacy protecting of ToSC. By training models in an adversarial setting, it is possible to generate feature representations that are informative for intended tasks but are simultaneously difficult for adversarial models to exploit. This is particularly useful in scenarios where sensitive data needs to be protected from potential privacy breaches. The adversarial process ensures that while the essential characteristics of the data are preserved for the task at hand, the ability of an adversary to extract sensitive information from the data is minimized. This approach is instrumental in developing models that balance the need for utility and privacy.
It is noteworthy that the transmitter does not know the architectures of the adversary neural networks in the design. In practice, the adversary is not fixed, and it is impractical to obtain the adversary’s network architecture. To defend against model inversion attacks, one can construct a simulated adversary and utilize adversarial learning to assist in ToSC training [11]. Even if the adversary’s inversion network architecture is unknown, the approach still provides effective privacy protection via a simulated adversary.
Besides adversary learning, LBVQ uses discrete latent representations which are more robust against inversion. This means that it’s harder to reconstruct the original input data accurately from the latent representations, thereby providing a layer of privacy. The quantization step in LBVQ, which converts continuous latent variables into discrete ones, acts as a bottleneck, reducing the amount of detailed information that can be decoded from the latent space. This inherent characteristic of LBVQ makes it a suitable choice for tasks where preserving privacy is crucial, as it inherently limits the detailed information that can be extracted from the encoded representations.
The integration of LBVQ within ToSC systems presents an additional benefit regarding compatibility with existing digital communication frameworks. In current ToSC models, features extracted by neural networks are typically continuous, aligning well with analog communications but not with prevalent digital communication systems. LBVQ, on the other hand, produces discrete representations that can be directly mapped to digital modulation symbols. This compatibility is instrumental in facilitating a seamless transition from well-established digital communication systems to the emerging ToSC paradigms, thereby bridging the gap between current technologies and future communication methodologies.
TABLE I: Complexity, cost and latency of four privacy-preserving mechanisms evaluated in the CIFAR-10 dataset.
@@ III-C Transceiver Structure Comparison
Comparing feature perturbation methods and encryption with intrinsic security incorporation strategies reveals distinct strengths and weaknesses. Their differences in transceiver structures are demonstrated in Fig. 2. Feature perturbation methods like differential privacy and encryption offer robust theoretical guarantees for privacy. Differential privacy provides a quantifiable measure of privacy by adding noise to the features. Encryption provides strong security for DeepJSCC output in transit but does not address the unique challenges of vulnerability to channel errors and real-time processing. On the other hand, adversarial learning offers an intrinsic security incorporation approach in ToSC. By introducing the adversary loss in the training phase of DeepJSCC, attacking-robust communication models can be obtained. Adversarial learning maintains strong performance even in the face of unknown or intentional attacks. However, the complexity and computational demands of adversarial models are notable challenges in the training phase. Additionally, continuous updates and model improvements may be needed to cope with new attack methods. LBVQ offers an effective approach for safeguarding feature privacy through the map of features into smaller vector spaces. This method significantly reduces feature dimensions while upholding the integrity of information. Nevertheless, the process of dimension reduction might result in potential information loss. And another challenge is the need for appropriate vector quantization based on the specific tasks and data types involved.
@@ III-D Cost and Delay Comparison
Tab. I shows the computational cost, complexity, and communication latency of four privacy schemes running on an 11th Gen Intel(R) processor at 2.50 GHz and a single 3060 CPU core, including the specific number of floating-point operations (FLOPs) in the whole training process, the number of parameters in all the neural network model, the time for a single training epoch (batch-size: 512) and the task-inference time of a single image input.
The mechanisms evaluated are DeepJSCC, DeepJSCC-DP, DeepJSCC-Encryption, IBAL, and DeepJSCC-LBVQ. The results indicate that DeepJSCC, DeepJSCC-DP, and DeepJSCC-Encryption demonstrate similar computational requirements and maintain a lower complexity profile, as evidenced by their FLOPs and parameter counts. In contrast, IBAL and DeepJSCC-LBVQ necessitate substantially greater resources, manifesting in increased FLOPs, a higher number of parameters, and extended training durations. Despite this, the inference times for all configurations, with the exception of DeepJSCC-LBVQ, remain comparably low. This suggests that all configurations, barring the LBVQ variant, achieve high efficiency during the inference phase. The LBVQ variant, however, incurs additional delays due to its discrete codebook mapping and remapping processes. Importantly, when compared to the baseline DeepJSCC model without privacy enhancements, the DeepJSCC-DP, DeepJSCC-Encryption, and IBAL models exhibit negligible increases in task-inference time, rendering them particularly suitable for applications requiring low-latency and privacy-sensitive remote inference.

### v2

# Semantic-Enabled 6G Communication: A Task-oriented and Privacy-preserving Perspective

@@ Abstract
Task-oriented semantic communication (ToSC) emerges as an innovative approach in the 6G landscape, characterized by the transmission of only vital information that is directly pertinent to a specific task. While ToSC offers an efficient mode of communication, it concurrently raises concerns regarding privacy, as sophisticated adversaries might possess the capability to reconstruct the original data from the transmitted features. This paper provides an in-depth analysis of privacy-preserving strategies specifically designed for ToSC relying on deep neural network-based joint source and channel coding (DeepJSCC). Our study encompasses a detailed comparative assessment of trustworthy feature perturbation methods such as differential privacy (DP) and encryption, alongside intrinsic security incorporation approaches like adversarial learning to train the JSCC and learning-based vector quantization (LBVQ). Our comparative analysis underscores the integration of advanced explainable learning algorithms into communication systems, positing a new benchmark for privacy standards in the forthcoming 6G era.
@@ I Introduction
The integration of artificial intelligence (AI) in 6G networks is anticipated to revolutionize industries by enabling new business models and services [1]. The network’s reliability, trustworthiness and timeliness will be critical in the scenario of massive mobile users with real-time response requirements [2]. Despite the promising prospects, the practical implementation continues to encounter numerous unprecedented hurdles, particularly for burst communications. Task-oriented semantic communication (ToSC), emerging as a promising paradigm, is primarily characterized by its selective transmission of information [3]. The fundamental concept of ToSC is to enhance communication efficiency and task performance by transmitting task-relevant semantic information, rather than raw data. This approach has garnered considerable attention, chiefly due to its proficiency in enhancing efficiency and reducing latency through the minimization of data transmission volume. For instance, a semantic communication approach with limited knowledge representation is proposed in [4], while a multimodal framework driven by AI models is introduced in [5] to achieve low-latency. Furthermore, ToSC exhibits the capacity to offer a more customized and efficient user experience. The targeted and efficient nature of ToSC, therefore, not only aligns with the technological advancements envisaged in 6G communications but also caters to the nuanced demands of modern digital applications, ensuring a seamless and user-oriented interaction.
@@
To alleviate the misunderstandings and incorrect interpretations, ToSC can increase information clarity, relevance, transparency, credibility, and verifiability by concentrating on task-relevant information, implementing well-designed mechanisms of channel coding and feedback. Although this strategy inherently provides a certain level of privacy since selective data transmission eliminates the unnecessary data sharing, the information conveyed may still be vulnerable [6]. If intercepted, even these task-relevant data bits could reveal personal or sensitive information [7]. This privacy-leakage risk is heightened by the advanced capabilities in machine learning (ML) and data analysis, which might enable adversaries to extract significant insights from minimal data. Compared to a generic ToSC framework [3] for various tasks with diverse data types, this paper has refined it for privacy and security applications. First, we discuss ongoing research aimed at developing privacy-preserving methods specifically for ToSC relying on deep neural network-based joint source and channel coding (DeepJSCC). Second, we aim to devise techniques that can be seamlessly incorporated into ToSC without significantly compromising their efficacy and efficiency, or in other words, strike a balance between utility, efficiency, and privacy. At last, we evaluate and contrast feature perturbation methodologies, such as differential privacy (DP) and encryption techniques, with intrinsic security incorporation approaches like adversarial learning and learning-based vector quantization (LBVQ). Our analysis highlights the potential for integrating sophisticated learning algorithms into contemporary communication systems.
@@ II Shift From Task-Agnostic Communications to ToSC and Privacy Challenges
Within the traditional framework of source-channel separation, the identification, representation, and transmission of information are rigorously addressed by rate-distortion theory and channel coding theory, respectively. This paradigm, which prioritizes reconstruction-oriented compression and task-agnostic communication, has underpinned several iterations of digital communication systems. However, the advent of machine-to-machine communications and human-machine interactions necessitates a reassessment of this paradigm, considering that exact reconstructions are often of secondary importance from a machine’s perspective. Notably, task-specific descriptors, derived via ML algorithms from latent feature spaces, are substantially more concise than their counterparts used for reconstruction purposes. Furthermore, communication systems trained end-to-end can significantly surpass those designed based on source-channel separation, across various performance metrics.
@@
In ToSC systems, data transmission is meticulously tailored to align with the receiver’s requirements. As shown in Fig. 1, different from traditional transceivers, the ToSC system comprises two main components: the “DeepJSCC” and the “Task inference module”. The DeepJSCC contains a JSCC encoder responsible for task-related information extraction, compression, and protection (against both channel noise and adverseries’ attacks). The output passes through a channel to the task inference module, which can handle tasks such as classification, object detection, information retrieval, and content generation. Inside, it displays a multi-layered network topology that indicates intricate processing is occurring to perform the given tasks.
@@
Despite these advantages, ToSC introduces significant privacy concerns. The main issue stems from the nature of the information being transmitted. Although ToSC systems transmit only task-specific information, this data can be sensitive and vulnerable to privacy infringements. Given the unpredictability of adversaries’ objectives, it is imperative to devise a comprehensive and efficacious strategy for safeguarding a spectrum of private data. For instance, in the transmission of facial images, adversaries may undertake diverse strategies to extract personal attributes such as gender or skin tone, or alternatively, engage in face recognition. In our research, we postulate the quality of image reconstruction—evaluated by metrics such as mutual information (MI) leakage—as a quantifiable metric for privacy preservation. Our approach assumes that amplifying distortions in data reconstructed by potential intruders indirectly shields various aspects of personal information. This approach resonates with the principles of perturbative privacy preservation, notably exemplified by DP paradigms. By escalating the degree of distortion, we can significantly diminish the likelihood or amplify the challenge for adversaries in gleaning sensitive information, thereby strengthening the robustness of privacy protection.
@@ III Privacy Preservation Methods for ToSC
Most ToSC systems are based on DeepJSCC architectures and employ an end-to-end training methodology to extract high-dimensional task-related channel-robust features for transmission. This approach ensures a coherent and automatically optimized process from data input to the final task output. Many traditional privacy protection techniques, such as $k$-anonymity, $l$-diversity, and $t$-closeness, are often designed to protect user privacy by modifying data and are not suitable for working with complex or high-dimensional data [7], as the process of making records indistinguishable can lead to significant data loss or impracticality in datasets with numerous attributes. A thorough analysis of privacy-preserving strategies specifically designed for ToSC is provided, along with a comparative assessment of feature perturbation methods and intrinsic security incorporation approaches.
@@ III-A Feature Perturbation Methods
Conventional approaches concentrate on the safeguarding of original data, typically through direct pre-processing under the assumption that the entire AI model is in the possession of a potentially untrustworthy third party. However, in DeepJSCC-based ToSC, the encoder belongs to the transmitter, which plays the role of task-related information extraction, compression, and protection. The employment of DeepJSCC makes conventional techniques applied in-between source and channel coding difficult in DeepJSCC. Applying these techniques will prevent the DeepJSCC to extract the task-related and channel-robust features, thus the only feasible is to process the output of DeepJSCC. Therefore, traditional data perturbations should be replaced by the feature perturbations on the DeepJSCC output. Next, we discuss two typical feature perturbation methods for privacy preservation.
  • DP provides a mathematical framework to quantify privacy loss and offers strong guarantees by adding noise to DeepJSCC outputs. Its robustness and theoretical assurances make it widely adopted. However, the noise added to protect individual data points can reduce data utility, particularly in scenarios requiring precise semantic information for task inference. While excessive noise enhances privacy, it inherently degrades task inference performance, impacting the utility of transmitted data.
  • Encryption serves as a fundamental method for securing DeepJSCC outputs by making them unintelligible without a decryption key. However, it can slow down training and inference. Without channel coding, encryption is also vulnerable, potentially compromising decryption integrity. To address these issues, simple or error-robust encryption methods, like key-guided feature shuffling [8], are recommended for ToSC. Additionally, encryption-based methods face key-sharing burdens and the key leakage.
@@ III-B Intrinsic Security Incorporation Strategies
Another approach involves designing DeepJSCC with intrinsic security. Two typical strategies are as follows.
  • Adversarial learning trains models in a competitive setup involving a generator and a discriminator. The generator creates data indistinguishable from real data, while the discriminator aims to differentiate between fake and real data, enhancing both models’ performance. This technique is vital for ToSC privacy protection. It generates feature representations useful for tasks yet hard for adversaries to exploit, crucial in protecting sensitive data. The adversarial process preserves data’s essential characteristics while minimizing an adversary’s ability to extract sensitive information, balancing utility and privacy. Notably, the transmitter is unaware of the adversary’s neural network architecture, which is impractical to obtain. A simulated adversary can be constructed and used in adversarial learning to assist ToSC training [9], defending against model inversion attacks (MIAs). MIAs, a common privacy threat to ML models, involve adversaries using access to trained models to reconstruct sensitive training data. Even without knowing the adversary’s inversion network architecture, this approach effectively protects privacy through a simulated adversary.
  • LBVQ uses discrete latent representations which are more robust against inversion. This means that it’s harder to reconstruct the original input data accurately from the latent representations. The quantization step in LBVQ, which converts continuous latent variables into discrete ones, acts as a bottleneck, reducing the amount of detailed information that can be decoded from the latent space. This inherent characteristic of LBVQ makes it a suitable choice for tasks where preserving privacy is crucial. The integration of LBVQ within ToSC systems presents an additional benefit regarding compatibility with existing digital communication frameworks. In current ToSC models, features extracted by neural networks are typically continuous, aligning well with analog communications but not with prevalent digital communication systems. Fortunately, LBVQ produces discrete representations that can be directly mapped to digital modulation symbols. This compatibility is instrumental in facilitating a seamless transition from well-established digital communication systems to the emerging ToSC paradigms.
TABLE I: Computation complexity, model training cost and learning latency of 4 privacy-preserving mechanisms over CIFAR-10 dataset.
@@ III-C Transceiver Structure Comparison
Comparing feature perturbation methods and encryption with intrinsic security incorporation strategies reveals distinct strengths and weaknesses, as demonstrated in Fig. 2. Feature perturbation methods like DP and encryption offer robust theoretical guarantees for privacy. DP provides a quantifiable measure of privacy by adding noise to the features. Encryption provides strong security for DeepJSCC output in transit but does not address the unique challenges of vulnerability to channel errors and real-time processing. While adversarial learning offers an intrinsic security incorporation approach in ToSC. By introducing the adversary loss in the training phase of DeepJSCC, attacking-robust communication models can be obtained. Adversarial learning maintains strong performance even in the face of unknown or intentional attacks. However, the complexity and computational demands of adversarial models are notable challenges in the training phase. Additionally, continuous updates and model improvements may be needed to cope with new attack methods. LBVQ offers an effective approach for safeguarding feature privacy through the map of features into smaller vector spaces. This method significantly reduces feature dimensions while upholding the integrity of information. Nevertheless, the process of dimension reduction might result in potential information loss. And another challenge is the need for appropriate vector quantization based on the specific tasks and data types involved.
@@ III-D Cost and Delay Comparison
Table I shows the computational cost, complexity, and communication latency of four privacy schemes running on an 11th Gen Intel(R) processor at 2.50 GHz and a single 3060 CPU core, including the specific number of floating-point operations (FLOPs) in the whole training process, the number of parameters in all the neural network model, the time for a single training epoch (batch-size: 512) and the task-inference time for a single image. The results indicate that DeepJSCC, DeepJSCC-DP, and DeepJSCC-Encryption demonstrate similar computational requirements and maintain a lower complexity profile, as evidenced by their FLOPs and parameter counts. In contrast, IBAL and DeepJSCC-LBVQ necessitate substantially greater resources, manifesting in increased FLOPs, a higher number of parameters, and extended training durations. Despite this, the inference times for all configurations, with the exception of DeepJSCC-LBVQ, remain comparably low. This suggests that all configurations, barring the LBVQ variant, achieve high efficiency during the inference phase. The LBVQ variant, however, incurs additional delays due to its discrete codebook mapping and remapping processes. Importantly, when compared to the baseline DeepJSCC model without privacy enhancements, the DeepJSCC-DP, DeepJSCC-Encryption, and IBAL models exhibit negligible increases in task-inference time, rendering them particularly suitable for applications requiring low-latency and privacy-sensitive remote inference.